# AL-QASIDA: Analyzing LLM Quality and Accuracy Systematically in Dialectal Arabic

### v1

@@ Abstract
Dialectal Arabic (DA) varieties are under-served by language technologies, particularly large language models (LLMs). This trend threatens to exacerbate existing social inequalities and limits language modeling applications, yet the research community lacks operationalized LLM performance measurements in DA. We present a method that comprehensively evaluates LLM fidelity, understanding, quality, and diglossia in modeling DA. We evaluate nine LLMs in eight DA varieties across these four dimensions and provide best practice recommendations. Our evaluation suggests that LLMs do not produce DA as well as they understand it, but does not suggest deterioration in quality when they do. Further analysis suggests that current post-training can degrade DA capabilities, that few-shot examples can overcome this and other LLM deficiencies, and that otherwise no measurable features of input text correlate well with LLM DA performance.
@@ 1 Introduction
Many LLMs today are proficient in MSA but reluctant to model DA; users wishing to converse in DA often have to perform extensive prompt acrobatics to coerce the LLM to use the dialect.333For instance, when conversing with ChatGPT in Egyptian Arabic, it took the author 14 turns of conversation before the LLM used the desired dialect (see Table 5). Because MSA proficiency correlates with socioeconomic advantage among Arabs, LLMs’ singular MSA aptitude may exacerbate existing inequalities. And even when interacting with MSA-proficient users, replying to informal DA inputs with formal MSA sounds unnatural and limits an LLM’s uses.
LLM pre-training data is diverse and likely includes content in many Arabic dialects. However, post-training uses existing human-labeled data, which is typically MSA, and newly collected data that commonly adopt an official tone corresponding to a formal MSA register. Current LLM behavior suggests that they have the capability to both understand and model DA, but that they default to MSA regardless, a behavior that could be caused by non-language-modeling objectives such as post-training. We thus hypothesize:
@@
AL-QASIDA: a method to evaluate LLM DA proficiency along dimensions of fidelity, understanding, quality, and diglossia
@@
Best-practice recommendations, including use of few-shot prompts for DA generation, Llama-3 for monolingual tasks, and GPT-4o for cross-lingual requests of Egyptian or Moroccan varieties
@@
Six key findings from our AL-QASIDA evaluation and analysis
Findings from our central evaluation include:
@@
Current post-training can bias LLMs against DA, but otherwise improves text quality.
@@ 3 Methodology
To be proficient in a DA, an LLM requires different competencies that we define as the following:
@@
Fidelity: Can the LLM identify and produce the correct DA variety when prompted to?
@@
Quality: Is the LLM able to model the DA variety well? (I.e. Does its quality deteriorate compared to MSA or another language?)
@@
Diglossia: Does the LLM understand how to translate between DA and MSA?
@@ 3.1 Operationalizing fidelity
Evaluating fidelity requires a NADI model to identify the Arabic variety of LLM outputs. In step with Abdul-Mageed et al. (2024), we note that single-label NADI classifications can lead to false errors in evaluation, like that seen in Figure 2. Accordingly, we apply the NADI 2024 shared task baseline model444Originally trained for the single-label task of Abdul-Mageed et al. (2023) Abdul-Mageed et al. (2024), but we extract the probability of the desired dialect from the output logits rather than a one-hot country-level classification. The NADI score () of an LLM is then the probability that its output is in the desired DA variety (denoted by country ). Because Keleg et al. (2023) and Abdul-Mageed et al. (2024) established NADI and ALDi as independent (see §2.1), this NADI model does not distinguish between MSA and DA: only between country-level varieties. Hence, since we want output (1) in the requested variety and (2) in DA (as opposed to MSA), our evaluation also requires an ALDi model. We define a DA fidelity performance metric based on both these desiderata: Arabic Dialect Identification And DIalectness (ADI2) score. Given LLM output ,555Because some LLMs, especially base models, tended to repeat DA inputs in monolingual settings, we remove any copies of the prompt from output before ADI2 computation. we define:
Because some DA varieties are close (see §2.1), we also compute an alternative to , which we call the NADI macro-score and define , where , each region itself representing a set of countries . Accordingly, the ADI2 macro-score , given . This metric indicates whether the LLM responds with DA varieties from the right general region, and thus allows for NADI confusions between proximate varieties.
@@ 3.2 Operationalizing other competencies
We use DA-to-English translation to evaluate the LLM’s understanding of DA text. As English is the closest language to an LLM’s internal representation Etxaniz et al. (2024), we can assess to what extent the model understood the original DA text based on the quality of its English translation. Machine translation (MT) metrics such as BLEU Papineni et al. (2002) or human annotations thus serve as proxies to DA understanding. We employ SpBLEU Goyal et al. (2022) and chrF Popović (2015). In addition to assessing understanding via MT, we ask human annotators to evaluate understanding directly by determining to what extent the LLM fulfilled requests in monolingual DA prompts during the fidelity evaluation. LLM-as-a-judge evaluators Zheng et al. (2023) are less accurate when rating responses in low-resource varieties such as DA, but our approach could be adapted to use such evaluators in the absence of human annotators.666optionally translating outputs to MSA before annotation At this stage we also ask annotators for judgments on fluency and dialectness, which we use in the next phase of evaluation.
@@
We next measure the LLM’s fluency and semantic accuracy in DA, first by English-to-DA MT. Again treating English as a proxy for the LLM’s semantic knowledge, this measures how well the LLM can express a variety of semantic concepts in DA. In this setting, we supplement automatic MT scores with human evaluations of adequacy and fluency of translations, as well as dialectness judgments, detailed in §4.2.777Again in the absence of human annotators, LLM-as-a-judge could be substituted at this stage. We couple this with correlative evaluation to reveal whether the LLM’s output quality deteriorates in DA compared to MSA. For this we gather all human (or LLM-as-a-judge) DA fluency annotations previously collected and correlate them with dialectness scores, to determine whether fluency degrades in more dialectal generation.
@@ 3.3 Evaluation corpora
For the cross-lingual corpus, we adapted a set from Marchisio et al. (2024) of English LLM prompts with explicit requests for responses in different languages, and substituted the names of DA varieties instead. This set originally came from three distinct collections of LLM user prompts: a subset of Okapi Lai et al. (2023) inputs to the Alpaca LLM Taori et al. (2023), a collection of ChatGPT inputs scraped from the ShareGPT API, and a corpus of human-curated prompts commissioned by Marchisio et al. (denoted Cohere). For the monolingual and bitext corpora, meanwhile, we selected four existing DA data sets based on their style diversity and dialectal coverage.
@@
We used two multi-variety DA bitext corpora, integrating 200 sentence pairs from each in our bitext corpus and 100 sentences from the monolingual DA portion of each in our monolingual corpus. The first, MADAR-26 Bouamor et al. (2018), is a multi-way parallel bitext with English, MSA, and DA from 25 Arab League cities. The English source sentences were sourced from the Basic Traveling Expression Corpus (BTEC) Takezawa et al. (2007) and manually translated into the 26 Arabic varieties.888We used sets for Riyadh, Damascus, Jerusalem, Khartoum, Cairo, Algiers, and Fes to render seven countries’ DA. The genre of this corpus is BTEC, i.e. everyday utterances that might be expressed verbally. The second set, FLORES-200 NLLB Team et al. (2022), is an MT evaluation benchmark of 1012 sentences in 204 language varieties. The English source texts were sampled from wiki sites and then translated manually. We use the sets for Najdi, North Levantine, South Levantine, Egyptian, and Moroccan Arabic to represent KSA, Syria, Palestine, Egypt, and Morocco, respectively.
@@
We then used two multi-variety monotext DA corpora, adding 100 additional sentences per variety from each to our monolingual corpus. The first, MADAR-Twitter Bouamor et al. (2019), contains 2,980 tweets from 4,375 profiles in 21 Arab countries. The tweets were seeded from 25 hashtags representing Arab League states (e.g. #Kuwait, #Egypt) and labeled by country. The second corpus, HABIBI El-Haj (2020), contains 30k song lyrics by artists from 18 Arab countries.999A native speaker of two DA varieties manually cleaned mislabeled sentences from the HABIBI sets we used, since some artists wrote in a DA variety other than their own. This annotator also deliberately selected lyrics for our 100-sentence subsets to highlight distinctions between varieties. We used eight country-level subsets from each of these collections (corresponding to all eight countries in Figure 1).
@@
Note that the purpose of monolingual evaluation is to measure how well an LLM matches a user’s input DA variety, so it should be composed of the commands instruction fine-tuned LLMs are accustomed to. The data sources we used for monolingual sentences provided generic sentences of four genres: BTEC, wiki text, tweets, and song lyrics. To transform these diverse generic sentences into instructions, we used eight instruction templates (shown in Table 7). We surveyed native speakers of the eight DA varieties covered to retrieve translations of the templates. Then for each generic sentence, we randomly selected one of the templates in the appropriate variety and inserted the generic sentence, transforming it into a command. Also randomly chosen along with the template was the sentence location: i.e. whether it was inserted at the start or end of the template, or in the middle where applicable (see Table 7).
@@ 4 Evaluation and Results
We detail our evaluation of nine LLMs for eight country-level DA varieties:101010listed roughly from east to west Kuwait, Saudi Arabia, Syria, Palestine, Sudan, Egypt, Algeria, and Morocco. These constitute two varieties from each of four Arabic dialectal regions (see Figure 1).
@@ 4.1 Primary results
Recall from §3 that we evaluate DA fidelity (ADI2 score and macro-score), and MT in four directions: DAEnglish and DAMSA. We begin with DA fidelity and present scores across eight dialects, four genres, and nine LLMs. Six are competitive open-source LLMs Command-R (and base model),111111https://cohere.com/blog/command-r Command-R+ (and base model),121212https://cohere.com/blog/command-r-plus-microsoft-azure and Llama 3.1 (and base model) Dubey et al. (2024). One is a competitive closed LLM, GPT-4o.131313https://openai.com/index/hello-gpt-4o/ And two are LLMs recognized for specialty in Arabic: ACEGPT Huang et al. (2024) (selected because of its prominence) and SILMA Team (2024) (selected because it led the Arabic LLM leaderboard141414https://huggingface.co/spaces/OALL/Open-Arabic-LLM-Leaderboard, as of 27-10-2024 during our evaluation). Results for monolingual and cross-lingual prompts are in Figures 3 and 4, respectively (more detailed results in Appendix A).
The order of model performance is roughly consistent across genres and dialects. In monolingual settings, the Command-R+ base model, Llama-3.1, and Llama-3.1 base model performed best,151515impressively, given its 8B parameters (see Table 9). followed by Command-R base. Command-R, Command-R+, GPT-4o, and ACEGPT typically perform worse, and SILMA worst of all (see Fig. 3). Almost all ADI2 scores fell below 50%. The Command base models’ higher performance supports Hypothesis 1 that post-training can inhibit DA modeling. All LLMs performed poorly on the cross-lingual task. Base models, ACEGPT, and Llama 3.1 were unable to complete the cross-lingual task whatsoever, and the performance of Command-R, Command-R+, and SILMA was low. GPT-4o was the only model to surpass 13% ADI2, and only on half of the dialects. We summarize results in Figure 4.161616See Figure 9 for fuller results. Note that on the lowest performing DA varieties, no LLMs exceed 20% ADI2 on any genre for either task, and that even on Egyptian and Moroccan, a majority of LLMs still score below this threshold.
@@ 4.2 Human evaluation
We asked two native speakers of Egyptian and one of Syrian Arabic to make judgments of the Command-R+ and GPT-4o DA outputs for both MT and monolingual tasks, as well as Command-R+ base for the monolingual task. Annotators judged Adherence of monolingual responses, or how well the LLM fulfilled the user’s request, and Adequacy of translations, or how well a translation reflected the meaning of the original sentence. For both tasks they made judgments of Arabic Fluency and Dialectal Fidelity, the final metric being the only one associated with the LLM’s ability to use the right DA variety. We defined the scales and guidelines for each of these measurements as shown in Figure 6 and Table 10. We averaged human scores across 50 prompts for each model. Annotators received prompts and completions in random order without LLM labels, for unbiased review.
See Figure 6 for average human eval scores. Results indicate that the instruction fine-tuned models excel at (1) producing fluent Arabic text, (2) translating into Arabic with semantic adequacy, and (3) fulfilling DA requests, but that they struggle to do so in the correct DA variety. Their high adequacy and low dialectal fidelity scores (along with low ADI2 scores and high DAEng MT scores in §4.1) indicate LLMs’ stronger DA understanding than generation, supporting Hypothesis 2. Command-R+ base has better dialectal fidelity, in part due to prompt reduplication, but fulfills user requests poorly with low fluency. We thus conclude that post-training can harm DA proficiency, but appears needed for other aspects of text quality.
@@ 5 Follow-up Experiments and Analysis
For feature-level correlations with ADI2 score, we performed Spearman’s rank tests for numerical features and ANOVA tests for categorical features to find and coefficients. We display these only for the relationships where , in Table 4, for both monolingual and cross-lingual tasks. No values in the table exceed Adams and Conway’s (2014) threshold of strong correlation, , and no values exceed magnitude . It seems target DA variety, LLM used, and number of few-shot examples correlate moderately with ADI2; but other features have weak or no correlation. In the monolingual task, the dialectness of the prompt also correlates significantly but weakly with both ADI2 and output dialectness ( and , respectively).171717High input dialectness can result in a range of output dialectness, but low input dialectness typically precludes high output dialectness. See Fig. 12. For a given LLM in a given DA variety, no feature of input text correlates strongly with DA performance.

### v2

@@ Abstract
Dialectal Arabic (DA) varieties are under-served by language technologies, particularly large language models (LLMs). This trend threatens to exacerbate existing social inequalities and limits LLM applications, yet the research community lacks operationalized performance measurements in DA. We present a framework that comprehensively assesses LLMs’ DA modeling capabilities across four dimensions: fidelity, understanding, quality, and diglossia. We evaluate nine LLMs in eight DA varieties and provide practical recommendations. Our evaluation suggests that LLMs do not produce DA as well as they understand it, not because their DA fluency is poor, but because they are reluctant to generate DA. Further analysis suggests that current post-training can contribute to bias against DA, that few-shot examples can overcome this deficiency, and that otherwise no measurable features of input text correlate well with LLM DA performance.
@@ 1 Introduction
Many LLMs today are proficient in MSA but reluctant to model DA; users wishing to converse in DA often have to perform extensive prompt acrobatics to coerce the LLM to use the dialect.333For instance, when conversing with ChatGPT in Egyptian Arabic, it took one of the authors 14 turns of conversation before the LLM used the desired dialect (see Table 5). Because MSA proficiency correlates with educational settings (and as a result socioeconomic advantage), such disparity may exacerbate existing inequalities. Even when interacting with MSA-proficient users, replying to informal DA inputs with formal MSA sounds unnatural and limits an LLM’s uses.
LLM pre-training data is diverse and likely includes content in many Arabic dialects. However, post-training uses existing human-labeled data, which is typically MSA, and newly collected data that commonly adopt an official tone corresponding to a formal MSA register. Current LLM behavior suggests that models have the capability to understand and model DA, but that they default to MSA regardless, which could be caused by biases introduced in post-training. We thus hypothesize:
@@
AL-QASIDA: a method to evaluate LLM DA proficiency along dimensions of fidelity, understanding, quality, and diglossia.
@@
Best practice recommendations, including use of few-shot prompts for DA generation, Llama-3 for monolingual tasks, and GPT-4o for cross-lingual requests of Egyptian or Moroccan varieties.
@@
The following six key findings from our AL-QASIDA evaluation and analysis.
These findings include:
@@
Post-training can bias LLMs against DA (see Hyp. 1), but otherwise improves text quality.
@@ 3 Methodology
To be proficient in DA, an LLM requires different competencies that we define as the following:
@@
Fidelity: Can the LLM identify and produce the correct DA variety when asked?
@@
Quality: Is the LLM able to model the DA variety well (i.e., does its quality deteriorate compared to MSA or another language)?
@@
Diglossia: Can the LLM translate between DA and MSA?
@@ 3.1 Operationalizing Fidelity
To this end, we require a NADI model to identify the Arabic variety of LLM outputs. As single-label NADI classifications can lead to false negatives in evaluation Abdul-Mageed et al. (2024), as seen in Figure 2. To mitigate this, we apply the NADI 2024 shared task baseline model444The model was originally trained for the single-label task of Abdul-Mageed et al. (2023). Abdul-Mageed et al. (2024), but extract the probability of the desired dialect from the output logits rather than using the one-hot classification output. The NADI score () of an LLM is then the probability that its output is in the desired DA variety of country . As NADI and ALDi are independent (see §2.1), this NADI model does not distinguish between MSA and DA: only between country-level varieties. We therefore employ an additional model for ALDi: to differentiate MSA and DA Keleg et al. (2023). We finally define a DA fidelity performance metric that combines both NADI and ALDi scores: Arabic Dialect Identification And DIalectness (ADI2) score. Given LLM output ,555Because some LLMs, especially base models, tend to repeat DA inputs in monolingual settings, we remove any copies of the prompt from output before ADI2 computation. we define:
Because some DA varieties are similar (see §2.1), we compute an alternative to , which we call the NADI macro-score: , where , with each region itself representing a set of countries . Accordingly, , given . This metric indicates whether the LLM responds with DA varieties from the correct region, and thus allows for NADI confusions between proximate varieties.
@@ 3.2 Operationalizing Other Competencies
We use DA-to-English translation to evaluate the LLM’s understanding of DA text. As English is the closest language to an LLM’s internal representation Etxaniz et al. (2024), we can assess to what extent the model understood the original DA text based on the quality of its English translation. Machine translation (MT) metrics such as BLEU Papineni et al. (2002) or human annotations thus serve as proxies to DA understanding. We employ SpBLEU Goyal et al. (2022) and chrF Popović (2015). In addition to assessing Understanding via MT, we ask human annotators to evaluate understanding directly by determining to what extent the LLM fulfilled requests in monolingual DA prompts during the Fidelity evaluation.666We rely on humans as LLM-as-a-judge evaluators Zheng et al. (2023) are less accurate when rating responses in low-resource varieties such as DA.
@@
We next measure LLMs’ fluency and semantic accuracy in DA, first by English-to-DA MT. Once more, we treat English as a proxy for the LLM’s semantic knowledge; this measures how well the LLM can express a variety of semantic concepts in DA. We supplement automatic MT scores with human evaluations of adequacy and fluency of translations, as well as dialectness judgments, detailed in §4.2. We couple this with correlative evaluation to reveal whether the LLM’s output quality deteriorates in DA compared to MSA. For this we gather all human DA fluency annotations previously collected and correlate them with dialectness scores, to determine whether fluency degrades in more dialectal generation. (See §4.2.)
@@ 3.3 Evaluation Corpora
For the cross-lingual corpus, we adapted a set from Marchisio et al. (2024) of English LLM prompts with explicit requests for responses in different languages, and substituted the names of DA varieties. This set originally came from three distinct collections of LLM user prompts: a subset of Okapi Lai et al. (2023) inputs to the Alpaca LLM Taori et al. (2023), a collection of ChatGPT inputs scraped from the ShareGPT API, and a corpus of human-curated prompts commissioned by Marchisio et al. (denoted Cohere). For the monolingual and bitext corpora, meanwhile, we selected four existing DA data sets based on their style diversity and dialectal coverage.
@@
We used two multi-variety DA bitext corpora, integrating 200 sentence pairs from each in our bitext corpus and 100 sentences from the monolingual DA portion of each in our monolingual corpus. The first, MADAR-26 Bouamor et al. (2018), is a multi-way parallel bitext with English, MSA, and DA from 25 Arab League cities. The English source sentences were sourced from the Basic Traveling Expression Corpus (BTEC) Takezawa et al. (2007) and manually translated into the 26 Arabic varieties.777We used sets for Riyadh, Damascus, Jerusalem, Khartoum, Cairo, Algiers, and Fes to render seven countries’ DA. The genre of this corpus is BTEC, i.e. everyday utterances that might be expressed verbally. The second set, FLORES-200 NLLB Team et al. (2022), is an MT evaluation benchmark of 1012 sentences in 204 language varieties. The English source texts were sampled from wiki sites and then translated manually. We use the sets for Najdi, North Levantine, South Levantine, Egyptian, and Moroccan Arabic to represent KSA, Syria, Palestine, Egypt, and Morocco, respectively.
@@
We then used two multi-variety monotext DA corpora, adding 100 additional sentences per variety from each to our monolingual corpus. The first, MADAR-Twitter Bouamor et al. (2019), contains 2,980 tweets from 4,375 profiles in 21 Arab countries. The tweets were seeded from 25 hashtags representing Arab League states (e.g. #Kuwait, #Egypt) and labeled by country. The second corpus, HABIBI El-Haj (2020), contains 30k song lyrics by artists from 18 Arab countries.888A native speaker of two DA varieties manually cleaned mislabeled sentences from the HABIBI sets we used, since some artists wrote in a DA variety other than their own. This annotator also deliberately selected lyrics for our 100-sentence subsets to highlight distinctions between varieties. We used eight country-level subsets from each of these collections (corresponding to all eight countries in Figure 1).
@@
Note that the purpose of monolingual evaluation is to measure how well an LLM matches a user’s input DA variety, so it should be composed of prompts instruction-tuned LLMs are accustomed to. The data sources we used for monolingual sentences provided generic sentences of four genres: BTEC, wiki text, tweets, and song lyrics. To transform these diverse generic sentences into instructions, we used eight instruction templates (shown in Table 7). We surveyed native speakers of the eight DA varieties covered to retrieve translations of the templates. Then for each generic sentence, we randomly selected one of the templates in the appropriate variety and inserted the generic sentence, transforming it into a command. Also randomly chosen along with the template was the sentence location: i.e. whether it was inserted at the start or end of the template, or in the middle where applicable (see Table 7).
@@ 4 Evaluation and Results
We detail our evaluation of nine LLMs for eight country-level DA varieties:999listed roughly from east to west Kuwait, Saudi Arabia, Syria, Palestine, Sudan, Egypt, Algeria, and Morocco. These constitute two varieties from each of four Arabic dialectal regions (see Figure 1).
@@ 4.1 Primary Results
Recall from §3 that we evaluate Fidelity, via ADI2 score and macro-score; and Understanding, Quality, and Diglossia, via DAEnglish and DAMSA MT. We begin with Fidelity and present scores across eight dialects, four genres, and nine LLMs including six open-source LLMs Command-R (and base model),101010https://cohere.com/blog/command-r Command-R+ (and base model),111111https://cohere.com/blog/command-r-plus-microsoft-azure and Llama 3.1 (and base model) Dubey et al. (2024); one closed LLM, GPT-4o;121212https://openai.com/index/hello-gpt-4o/ and two LLMs specialized in Arabic, ACEGPT Huang et al. (2024) (selected because of its prominence) and SILMA Team (2024) (selected because it led the Arabic LLM leaderboard131313https://huggingface.co/spaces/OALL/Open-Arabic-LLM-Leaderboard, as of 27-10-2024 during our evaluation). Results for monolingual and cross-lingual prompts are in Figures 3 and 4, respectively (more detailed results in Appendix A).
Regarding Fidelity, the order of model performance is roughly consistent across genres and dialects. In monolingual settings, the Command-R+ base model, Llama-3.1, and Llama-3.1 base model performed best,141414impressively, given its 8B parameters (see Table 9). followed by Command-R base. Command-R, Command-R+, GPT-4o, and ACEGPT typically perform worse, and SILMA worst of all (see Fig. 3). Almost all ADI2 scores fell below 50%. The Command base models’ higher performance supports Hypothesis 1 that post-training can inhibit DA modeling. All LLMs performed poorly on the cross-lingual task. Base models, ACEGPT, and Llama 3.1 were unable to complete the cross-lingual task whatsoever, and the performance of Command-R, Command-R+, and SILMA was low. GPT-4o was the only model to surpass 13% ADI2, and only on half of the dialects. We summarize results in Figure 4.151515See Figure 10 for fuller results. Note that on the lowest performing DA varieties, no LLMs exceed 20% ADI2 on any genre for either task, and that even on Egyptian and Moroccan, a majority of LLMs still score below this threshold.
@@ 4.2 Human evaluation
We asked two native speakers of Egyptian and one of Syrian Arabic to make judgments of the Command-R+ and GPT-4o DA outputs for both MT and monolingual tasks, as well as Command-R+ base for the monolingual task. Annotators judged Adherence of monolingual responses, or how well the LLM fulfilled the user’s request, and Adequacy of translations, or how well a translation reflected the meaning of the original sentence. For both tasks they made judgments of Arabic Fluency and Dialectal Fidelity, the final metric being the only one associated with the LLM’s ability to use the right DA variety. We defined the scales and guidelines for each of these measurements as shown in Figure 6 and Table 10. We averaged human scores across 50 prompts for each model. Annotators received prompts and completions in random order, for an unbiased review.
See Figure 6 for average human eval scores. Results indicate that the instruction fine-tuned models excel at (1) producing fluent Arabic text, (2) translating into Arabic with semantic adequacy, and (3) fulfilling DA requests, but that they struggle to do so in the correct DA variety. Their high adequacy and low dialectal fidelity scores (along with low ADI2 scores and high DAEng MT scores in §4.1) indicate LLMs’ stronger DA understanding than generation, supporting Hypothesis 2. This is an apparent reversal of the Generative AI Paradox West et al. (2024), which highlights that LLMs’ generative capabilities exceed their ability to understand their outputs. In contrast, we observe that LLMs are capable of understanding DA utterances but unable to produce fluent DA outputs.
Comparing base and post-trained models, Command-R+ base demonstrates better dialectal fidelity, in part due to prompt reduplication, but fulfills user requests poorly with low fluency. We thus conclude that post-training can harm DA proficiency, but is needed for other aspects of text quality such as instruction following ability.
@@ 5 Follow-up Experiments and Analysis
For feature-level correlations with ADI2 score, we performed Spearman’s rank tests for numerical features and ANOVA tests for categorical features to find and coefficients. We display these only for the relationships where , in Table 4, for both monolingual and cross-lingual tasks. No values in the table exceed Adams and Conway’s (2014) threshold of strong correlation, , and no values exceed magnitude . It seems target DA variety, LLM used, and number of few-shot examples correlate moderately with ADI2; but other features have weak or no correlation. In the monolingual task, the dialectness of the prompt also correlates significantly but weakly with both ADI2 and output dialectness ( and , respectively).161616High input dialectness can result in a range of output dialectness, but low input dialectness typically precludes high output dialectness. See Fig. 9. For a given LLM in a given DA variety, no feature of input text correlates strongly with DA performance.